Solve (z+1)^2 = 2z^2 for z.
Answer by Boreal(15235)

z^2+2z+1=2z^2
z^2-2z-1=0
z=(1/2)(2+/- sqrt(8))
=1+/- sqrt(2)
-
Check
this is 2+sqrt(2) squared or 4+4 sqrt(2)+2 or 6+4 sqrt(2)=2(2+2 sqrt(2)^2=4+4 sqrt(2)+2, which works
or 2-sqrt(2) squared or 4-4 sqrt(2)+2 or 6-4 sqrt(2)=2(1-2sqrt(2)+2)=2-4sqrt(2)+4, which also works.